OpenStudy (anonymous):

MEDAL_______Convert the mixed number 8 5/11 to an improper fraction

OpenStudy (anonymous):

Multiply the whole number by the denominator, what do you get?

OpenStudy (anonymous):

8 * 11

OpenStudy (anonymous):

88

OpenStudy (anonymous):

Add the answer from the last question to the numerator.

OpenStudy (anonymous):

88 + 5

OpenStudy (anonymous):

93

OpenStudy (anonymous):

Great, the answer to the last question is your new numerator and the denominator remains the same.

OpenStudy (anonymous):

\[\frac{ 93 }{ 11 }\]

OpenStudy (anonymous):

Yeah sure, post all of the question and I will get to them.

OpenStudy (anonymous):

Expresss the 3/6 in the simplest form * 1/2 Add 3/5 + 4/5= * 7 Subtract 6/6-4/6= * 2 Write two fractions equivalent to 4/5 * 8/10 12/15 Convert the mixed number 8 5/11 to an improper fraction * 93/11 Convert the decimal 2.7 to a fraction or mixed number Fraction 27/10 Mixed number 7/10 5/6 + 7/8 + 3 1/6= 4.875 What is the reciprocal of 8/3 * 3/8 Divide 2/3 by 1/3 * 2

OpenStudy (anonymous):

Express the 3/6 in the simplest form * 1/2 Correct! Add 3/5 + 4/5= * 7 Incorrect. Subtract 6/6-4/6= * 2 I will ask a question about this question after I post this, wait a moment, I am confused. Write two fractions equivalent to 4/5 * 8/10 12/15 Correct! Convert the mixed number 8 5/11 to an improper fraction * 93/11 Correct! (I hope) Convert the decimal 2.7 to a fraction or mixed number Fraction 27/10 Mixed number 7/10 The fraction is correct, the mixed number is not. 5/6 + 7/8 + 3 1/6= 4.875 Correct, however you want to convert it into a fraction. What is the reciprocal of 8/3 * 3/8 Correct! Divide 2/3 by 1/3 * 2 Yeah, about. Correct!

OpenStudy (anonymous):

5/6 + 7/8 + 3 1/6= 4.875 Convert the decimal 2.7 to a fraction or mixed number Fraction 27/10 Mixed number 7/10 The fraction is correct, the mixed number is not. Add 3/5 + 4/5= * 7 Incorrect. Subtract 6/6-4/6= * 2 I will ask a question about this question after I post this, wait a moment, I am confused.

OpenStudy (anonymous):

I shall also help out on those you got wrong.

OpenStudy (anonymous):

The first one you got wrong: Add 3/5 + 4/5= * 7

OpenStudy (anonymous):

It is not 7.

OpenStudy (anonymous):

That one may be a typo.

OpenStudy (anonymous):

When adding fractions of the same denominator, all you do is add the numerators.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

3 + 4 = 7

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

The answer for that one is 7/5

OpenStudy (anonymous):

As for the third question: 6/6 - 4/6 = 2/6 Only the numerator when the denominators are the same.

OpenStudy (anonymous):

Convert the decimal 2.7 to a fraction or mixed number Fraction 27/10 Mixed number 7/10

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

Your fraction is correct, however the mixed number is not.

OpenStudy (anonymous):

Mixed numbers are whole numbers with a fraction.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

The whole number for that was 2, then there is the .7, which can be 7/10 So: Convert the decimal 2.7 to a fraction or mixed number Fraction 27/10 Mixed number 2 7/10

OpenStudy (anonymous):

5/6 + 7/8 + 3 1/6= 4.875 Your response is correct, however your teacher may want it as a fraction.

OpenStudy (anonymous):

.875 is 7/8 as a fraction.

OpenStudy (anonymous):

So, 4 7/8

OpenStudy (anonymous):

XD

OpenStudy (anonymous):

Yeah, just doing one thing, I am preparing your new answer sheet.

OpenStudy (anonymous):

Express the 3/6 in the simplest form * 1/2 Add 3/5 + 4/5= * 7/5 Subtract 6/6 - 4/6= * 2/6 Write two fractions equivalent to 4/5 * 8/10 12/15 Convert the mixed number 8 5/11 to an improper fraction * 93/11 Convert the decimal 2.7 to a fraction or mixed number Fraction 27/10 Mixed number 2 7/10 5/6 + 7/8 + 3 1/6= 4 7/8 What is the reciprocal of 8/3 * 3/8 Divide 2/3 by 1/3 * 2

OpenStudy (anonymous):

I hope there is no error in that.

OpenStudy (anonymous):

Is there a question missing, or repeating?

OpenStudy (anonymous):

Kk. Have a nice day.

OpenStudy (anonymous):

I am fairly certain so. 2.7 = 27/10, or 2 and 7/10